OpenStudy (anonymous):

2. A calibration curve requires the preparation of a set of known concentrations of CV, which are usually prepared by dieting a stock solution whose concentration is known. Describe how to prepare 10. mL of a 5, 10, 15, and 20 µM CV solution using a 25 µM CV stock solution.

OpenStudy (cuanchi):

This is a typical dilution problem. Usually you can solve applying the formula \[C _{1}V _{1}= C _{2}V _{2}\] where C1 is the concentration of your stock solution , V1 the volume of this stock solution that you have to use to make the dilution, C2 the final concentration of your dilution and V2 the final volume of your diluted solution. For the 15 uM solution you will have Given: C1= 25uM, C2 = 20 uM , V2= 10 mL Find: V1 V1= C2V2/C1 = 15uM 10mL/20uM = 7.5mL To made the 20uM solution you have to put 7.5mL of your stock solution (25uM) and complete to 10 mL with (10-7.5= 2.5mL) of water For the 10uM solution you can use a similar calculation or figure out that you dilute in half the stock solution. Then 5mL of stock and 5ml of water. V1= C2V2/C1 = 10uM 10mL/20uM = 5mL For the 5uM also you can use the same formula or figure out this is 1/5 of the original solution you will need 2.5 mL of stock plus 7.5 of water V1= C2V2/C1 = 5uM 10mL/20uM = 2.5mL